OpenStudy (anonymous):

Calc 3. Set up, but do not evaluate, an iterated integral for the volume of the solid. Under the graph of f(x,y)=25-x^2-y^2 and above the xy-plane. I think the outer integral is 0 to 5, but I'm stuck on the inner integral because I can get it to depend on two variables, but not 1.

OpenStudy (anonymous):

Evaluate 5 and 3.14 pie and subtract. That and multiply!

OpenStudy (tkhunny):

In the x-y plane, z = 0, giving the circle x^2 + y^2 = 25. This is wonderfully symmetric, so work in just the first octant: \(4\cdot\int\limits_{0}^{??}\int\limits_{0}^{??} 25 - x^2 - y^2 \;dx\;dy\)